A man is 30 years older than his son. In 17 years his son would be half his age. How old is his son?
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement:
:
A man is 30 years older than his son.
M = S + 30
:
"In 17 years his son would be half his age."
(S+17) = .5(M + 17)
S + 17 = .5M + 8.5
S = .5M + 8.5 - 17
S = .5M - 8.5
:
How old is his son?
Substitute (S+30) for M in the above equation and do some basic algebra:
S = .5(S+30) - 8.5
S = .5s + 15 - 8.5
S - .5S = 6.5
.5S = 6.5
S = 13 yrs old is the son now
Then
M = 13 + 30 = 43 yrs old is the old man now
:
:
Check our solutions in the statement:
"In 17 years his son would be half his age."
(13 + 17) = .5(43 + 17)
30 = .5(60); confirms our solutions
:
:
Did I make this understandable to you? Any questions?